find the equation of the line through the point (5,6) which forma with the axes a triangle whose area is 80.


1
Expert's answer
2022-01-16T12:39:11-0500

Let the equation of the line is "y=mx+b, m\not=0."

"y" -intercept: "x=0, y=b"

Point "(0, b)."

"x" -intercept: "y=0, 0=mx+b=>x=-\dfrac{b}{m}"

Point "(0, b)."

The area of the triangle


"Area=\dfrac{1}{2}|b|\cdot|-\dfrac{b}{m}|=80"

"\dfrac{b^2}{160}=|m|"

Line passes through the point "(5, 6)"


"6=m(5)+b"

"m=\dfrac{6-b}{5}"

"b<6"

"\dfrac{b^2}{160}=\dfrac{6-b}{5}"

"b^2+32b-192=0"

"(b+16)^2=448"

"b=-16\pm8\sqrt{7}"

"b=-16-8\sqrt{7}"


"m=\dfrac{6-(-16-8\sqrt{7})}{5}=\dfrac{22+8\sqrt{7}}{5}"

"y=\dfrac{22+8\sqrt{7}}{5}x-(16+8\sqrt{7})"



"b=-16+8\sqrt{7}<6"


"m=\dfrac{6-(-16+8\sqrt{7})}{5}=\dfrac{22-8\sqrt{7}}{5}"

"y=\dfrac{22-8\sqrt{7}}{5}x+(-16+8\sqrt{7})"


"b>6"

"\dfrac{b^2}{160}=-\dfrac{6-b}{5}"

"b^2-32b+192=0"

"(b-16)^2=64"

"b=16\pm8"

"b=16-8=8"

"m=\dfrac{6-8}{5}=-\dfrac{2}{5}"

"y=-\dfrac{2}{5}x+8"

"b=16+8=24"

"m=\dfrac{6-24}{5}=-\dfrac{18}{5}"

"y=-\dfrac{18}{5}x+24"


Answer:


"y=\dfrac{22+8\sqrt{7}}{5}x-(16+8\sqrt{7})"

"y=\dfrac{22-8\sqrt{7}}{5}x+(-16+8\sqrt{7})"

"y=-\dfrac{2}{5}x+8"

"y=-\dfrac{18}{5}x+24"